OpenStudy (anonymous):

I need help with this negative 1 over 2 x + 3 = -x + 7?

OpenStudy (anonymous):



OpenStudy (akshay_budhkar):

-x/2+3=-x+7?

OpenStudy (akshay_budhkar):

if it is so then we have x-x/2=7-3 therefore x/2=4 therefore x=8

OpenStudy (akshay_budhkar):

r u getting it?

OpenStudy (anonymous):

Yeah., Thank you.

OpenStudy (akshay_budhkar):

np :D